OpenStudy (anonymous):

Solve this system of equations using the addition method. 4x - y = 1 3x + y = 13

OpenStudy (ivancsc1996):

well you just need to add the x's the y's and the results. You get: \[7x=14 \rightarrow x=2\]The you just solve for y on one of the equations:\[4x+y=1=4(2)+y=1 \rightarrow y=-7\]